# Long-Term Fairness Inquiries and Pursuits in Machine Learning: A Survey of Notions, Methods, and Challenges

### v2

@@ 7 Open Challenges and Future Works
Evaluating long-term and short-term fairness tension. Evaluating the tension and developing techniques for achieving fairness immediately and maintaining it over an extended period presents a significant challenge. Additionally, while the fairness-utility trade-off has been thoroughly examined in fairness literature (Mehrabi et al., 2021), assessing this long-term trade-off and establishing benchmarks to evaluate different methods remain important directions.
@@
Complexity of real-world dynamics. As we have seen in Sections 2.3 and 5, long-term fairness necessitates modeling and characterization of dynamics. However, the complexity of real-world scenarios, including limited access to demographic data, presents significant challenges for practitioners to fully understand the interplay of various dynamics and effectively address the discrimination involved. Translating insights from studies focused on specific applications to broader contexts and bridging the gap between empirical fairness evaluations (Gohar et al., 2023) and theoretical guarantees (Tang & Zhang, 2022) are important and ongoing challenges in the field.
@@
Intersectional identities. Current long-term fairness research has primarily concentrated on single-attribute group identities, such as gender or age. However, numerous studies reveal that bias is a nuanced phenomenon that spans multiple protected groups and intersecting identities Gohar & Cheng (2023). This complexity poses unique challenges, notably the exponential number of subgroups. Future endeavors should focus on developing methodologies and frameworks that capture and mitigate bias at the intersection of various attributes in long-term settings.
@@
Explainability over time. Explanability in AI has gained tremendous attention due to the need for transparency, which significantly affects acceptance within the broader community Gross (2007); Shin (2021). Long-term fairness notions, such as effort-based fairness, necessitate using explainable methods. It is vital to empower the impacted population to understand how decisions are made and determine strategies to enhance their qualifications in subsequent instances. Furthermore, incorporating explainability methods in dynamic fairness-aware recommendation environments is another direction Chen et al. (2019b).
@@
Long-term dynamics in other AI domains. Studies have shown that prompt engineering and fine-tuning in LLMs directly influence the model toward generating the desired outputs Zamfirescu-Pereira et al. (2023); Logan IV et al. (2021). In real-world scenarios, user interactions are dynamic and may even be strategic, guiding the LLMs in tailoring responses to their needs through repeated feedback. An important research direction is developing evaluation methods that capture these dynamics and assess bias and fairness in LLM (and other large models) outputs. On the other hand, it is also crucial to steer the response of LLMs to consistently align with human values represented by different demographics and cultures (Solaiman et al., 2023) in multi-round dialogues.
@@
Alignment with policy and regulatory framework in the long run. Trustworthy AI has drawn increasing attention from policy and regulatory efforts. For example, the European Commission High-Level Expert Group on AI presented the Ethical Guidelines (European Commission, 2019) that put forward a set of 7 key requirements an AI system should meet to be deemed trustworthy. The Organisation for Economic Cooperation and Development also proposed a framework to compare implementation tools for promoting Trustworthy AI (Organisation for Economic Co-operation and Development (2020), OECD). Consequently, it is important to consider long-term fairness together with other aspects of Trustworthy AI.
@@ 8 Conclusion
Recent research highlights the inadequacy of static one-shot fairness approaches in capturing unfairness over extended periods. To offer researchers an understanding of the field, we present a unifying review of long-term fairness in multiple problem settings and frameworks. First, we provide readers with a high-level overview of the conceptual frameworks employed to model long-term fairness problems. Next, we propose the first taxonomy on long-term fairness, wherein we explored long-term fairness definitions, mitigation methods, and representative evaluation techniques across various applications. We also discuss briefly other problem settings that tackle this issue. Finally, we discuss open challenges and opportunities of long-term fairness research. Ultimately, we hope our survey serves as a foundational resource for those new to the field, fostering an understanding of the complexities associated with long-term fairness and drawing attention to the potential societal consequences that may arise without effective long-term fairness measures.
@@ A.1 Survey Methodology
We employed several established approaches to conduct a comprehensive search for relevant papers under the umbrella of long-term fairness. Initially, we curated a set of seed papers that have discussed and evaluated long-term fairness problems (Zhang & Liu, 2021; Liu et al., 2018; Hu & Chen, 2018; D’Amour et al., 2020; Zhang & Liu, 2021; Zhang et al., 2020). Following common practice (Desolda et al., 2021; Snyder, 2019), we used these papers and domain knowledge to identify and develop a set of keywords to search for relevant work: “long-term fairness," “dynamic fairness," “(non-)static fairness," and “sequential fairness." Primarily, we sear‘ched proceedings of relevant conferences and journals that were likely to publish work that fit the definition of long-term fairness. These included but were not limited to FAccT, NeurIPS, AAAI, ICML, AIES, ICLR, and IJCAI. We also followed citation trails to and from these widely cited papers and querying across Google Scholar and Arxiv to identify additional related work.

### v3

@@ 8 Open Challenges and Future Works
The tension between long-term and short-term fairness.
@@
Evaluating the tension and developing techniques for achieving fairness immediately and maintaining it over an extended period presents a significant challenge. Approaches strictly oriented toward long-term goals may introduce short-term drawbacks, including reduced utility and fairness. Conversely, focusing predominantly on short-term fairness risks long-term fairness outcomes. Future work should investigate the trade-off between short-term and long-term fairness. Additionally, while the fairness-utility trade-off has been thoroughly examined in static fairness literature (Mehrabi et al., 2021), assessing this long-term trade-off and establishing benchmarks to evaluate different methods remain important directions.
Complexity of real-world dynamics.
@@
As we have seen in Sections 2.3 and 5, long-term fairness necessitates modeling and characterization of dynamics. However, the complexity of real-world scenarios, including limited access to demographic data, presents significant challenges for practitioners to fully understand the interplay of various dynamics and effectively address the discrimination involved. Translating insights from studies focused on specific applications to broader contexts and bridging the gap between empirical fairness evaluations (Gohar et al., 2023) and theoretical guarantees (Tang & Zhang, 2022) are important and ongoing challenges in the field.
Intersectional identities.
@@
Current long-term fairness research has primarily concentrated on single-attribute group identities, such as gender or age. However, numerous studies reveal that bias is a nuanced phenomenon that spans multiple protected groups and intersecting identities (Gohar & Cheng, 2023). This complexity poses unique challenges, notably the exponential number of subgroups. Future work should focus on developing methodologies and frameworks that capture and mitigate bias at the intersection of various attributes in long-term settings.
Explainability over time.
@@
Explainability in AI has gained tremendous attention due to the need for transparency, which significantly affects acceptance within the broader community (Gross, 2007; Shin, 2021). Long-term fairness notions, such as effort-based fairness, necessitate using explainable methods. It is vital to empower the impacted population to understand how decisions are made and determine strategies to enhance their qualifications in subsequent instances. Furthermore, incorporating explainability methods in dynamic fairness-aware recommendation environments is another direction (Chen et al., 2019b).
Long-term dynamics in other AI domains.
@@
Studies have shown that prompt engineering and fine-tuning in LLMs directly influence the model toward generating the desired outputs (Zamfirescu-Pereira et al., 2023; Logan IV et al., 2021). In real-world scenarios, user interactions are dynamic and may even be strategic, guiding the LLMs in tailoring responses to their needs through repeated feedback. An important research direction is developing evaluation methods that capture these dynamics and assess bias and fairness in LLM (and other large models) outputs. This, coupled with the interaction context, significantly influences LLM’s output (Zhang et al., 2021). On the other hand, it is also crucial to steer the response of LLMs to consistently align with human values represented by different demographics and cultures (Solaiman et al., 2023) in multi-round dialogues.
Alignment with policy and regulatory framework in the long run.
@@
Trustworthy AI has drawn increasing attention from policy and regulatory efforts. For example, the European Commission High-Level Expert Group on AI presented the Ethical Guidelines (European Commission, 2019) that put forward a set of 7 key requirements an AI system should meet to be deemed trustworthy. The Organisation for Economic Cooperation and Development also proposed a framework to compare implementation tools for promoting Trustworthy AI (Organisation for Economic Co-operation and Development (2020), OECD). Consequently, it is important to consider long-term fairness together with other aspects of Trustworthy AI.
@@ 9 Conclusion
Recent research highlights the inadequacy of static one-shot fairness approaches in capturing unfairness over extended periods. To offer researchers an understanding of the field, we present a unifying review of long-term fairness in multiple problem settings and frameworks. First, we provide readers with a high-level overview of the conceptual frameworks employed to model long-term fairness problems. Next, we propose the first taxonomy on long-term fairness, wherein we explored long-term fairness definitions, mitigation methods, and representative evaluation techniques across various applications. We also discuss briefly other problem settings that tackle this issue. Finally, we critically dissect the literature and discuss open challenges and opportunities of long-term fairness research. Ultimately, we hope our survey serves as a foundational resource for those new to the field, fostering an understanding of the complexities associated with long-term fairness and drawing attention to the potential societal consequences that may arise without effective long-term fairness measures.
Acknowledgments
Lu Cheng is supported by the National Science Foundation (NSF) Grant #2312862, NSF CAREER #2440542, NSF-Simons SkAI Institute, National Institutes of Health (NIH) #R01AG091762, and a Cisco gift grant. We would also like to acknowledge the support from NSF Award No. 2229881, AI Institute for Societal Decision Making (AI-SDM), the NIH under Contract R01HL159805, and grants from Quris AI, Florin Court Capital, and MBZUAI-WIS Joint Program. Zeyu Tang is supported by the National Institute of Justice (NIJ) Graduate Research Fellowship, Award No. 15PNIJ-24-GG-01565-RESS. We would also like to thank Nina Mehrabi for feedback on an earlier draft.
@@ A.1 Survey Methodology
We employed several established approaches to conduct a comprehensive search for relevant papers under the umbrella of long-term fairness. Initially, we curated a set of seed papers that have discussed and evaluated long-term fairness problems (Zhang & Liu, 2021; Liu et al., 2018; Hu & Chen, 2018; D’Amour et al., 2020; Zhang & Liu, 2021; Zhang et al., 2020). Following common practice (Desolda et al., 2021; Snyder, 2019), we used these papers and domain knowledge to identify and develop a set of keywords to search for relevant work: “long-term fairness,” “dynamic fairness,” “(non-)static fairness,” and “sequential fairness.” Primarily, we searched proceedings of relevant conferences and journals that were likely to publish work that fit the definition of long-term fairness. These included but were not limited to FAccT, NeurIPS, AAAI, ICML, AIES, ICLR, and IJCAI. We also followed citation trails to and from these widely cited papers and querying across Google Scholar and Arxiv to identify additional related work.